OpenStudy (anonymous):

Can someone how to do 3^3/2 by hand?

OpenStudy (anonymous):

explain**

OpenStudy (anonymous):

or any x^(y/z)

OpenStudy (unklerhaukus):

is that \[3^{3/2}\] or\[(3^3)/2\]

OpenStudy (anonymous):

the first one.. 3^(3/2)

OpenStudy (campbell_st):

so its really (3^3)^1/2 so do the inner part 1st (27)1/2 \[\sqrt{27}\] \[\sqrt{27}=\sqrt{9 \times3} =\sqrt{3^2 \times 3} = 3\sqrt{?}\]

OpenStudy (campbell_st):

oops ? = 3

OpenStudy (anonymous):

3sqrt3?

OpenStudy (sasogeek):

\[\ 3^\frac{3}{2}\sqrt{3^3}=\sqrt{9}=3 \]

OpenStudy (anonymous):

3^ 3/2 = surd 3^3 used, indices law.. surd 3^3 = surd 27 \[\sqrt{27}\] = \[\sqrt{(9)(3)}\] = \[3\sqrt{3}\]

OpenStudy (anonymous):

So what if it was 3^5/3?

OpenStudy (sasogeek):

\[\ \sqrt{27}=\sqrt{\text{9 x 3}} = \sqrt{9} \text{ x } \sqrt{3}= 9\sqrt{3} \]

OpenStudy (sasogeek):

I'm so slow at latex :(

OpenStudy (sasogeek):

\[\huge \sqrt[n]{A^b}=A^\frac{b}{n}\]

OpenStudy (anonymous):

for 3^(5/3) = \[\sqrt[3]{3^5}\] = \[\sqrt[3]{243}\] = \[\sqrt[3]{(27)(9)}\] = \[3 \sqrt[3]{9}\]

OpenStudy (sasogeek):

\[\huge note \ that \ \sqrt[2]{A}=\sqrt{A} \]